xXMarcelieXx:

someone help me with math proofs. how do you know which side is easier to work with for logical equivalence? @hero

xXMarcelieXx:

1 attachment
jhonyy9:

i think that you are right with this signed yelow but you know it why this is ?

xXMarcelieXx:

thats the given equation. i have no idea how to solve it

imqwerty:

it suffices to show that if \(\left[ \left( P \Lambda Q\right) \rightarrow R\right] , ~then~ \left(P \rightarrow R\right) V \left(Q \rightarrow R\right) \)

imqwerty:

Assume that \(\left[ \left( P \Lambda Q\right) \rightarrow R\right] \)

imqwerty:

that's how you start

imqwerty:

checkout some proofs/examples related to discreet mathematics